OpenStudy (anonymous):

What is the product of the square root of r(8 the square root of 2 -7)? Simplify.

OpenStudy (anonymous):

Possible Answers: a. 2t b. the square root of the quantity 2 times t c. 8 the square root of the quantity 2 times t -7 d. 8 the square root of the quantity 2 times t -7 the square root of t

OpenStudy (imstuck):

Is there a way you could post this expression in the equation editor so it makes a little more sense?

OpenStudy (anonymous):

Would it be this: \[\sqrt{t}(8\sqrt{2-7})\]

OpenStudy (anonymous):

The way I wrote it originally is exactly how it is written. I am having issues putting it into equations form. I think what I just posted is correct? Does it make sense?

OpenStudy (anonymous):

I think I got it \[8\sqrt{2t}-7\sqrt{t}\]

OpenStudy (imstuck):

Well, if it's that second expression you posted, it cannot be simplified because you cannot combine the square root of 2t and the square root of t because they are not like. And you cannot manipulate them to make them like. I think it might be this:\[t(8\sqrt{2}-7)\]If that be the case, then it would be 8sqrt 2 times t - 7t. Or \[8t \sqrt{2}-7t\]. Unless I know exactly what the expression is asking for, that is my best.